What is the electric flux through a cylindrical Gaussian surface at distance r for a wire of l and linear charge density given by d=kx

Grade:12

1 Answers

MASKED MAN
5 years ago
Given d=kx
And we know that electric field E of cylindrical Gaussian surface, E=k*2d
Then E=k*k*2x/x=2k*k